Find the inverse of the matrix
1 −1 4
2 9 8
6 5 2
using LU decomposition method.
1
Expert's answer
2020-04-20T16:48:54-0400

The LU decomposition "A = L U" of the given matrix is:

"\begin{pmatrix}
1 & -1 & 4 \\
2 & 9 & 8 \\
6 & 5 & 2
\end{pmatrix} = 
\begin{pmatrix}
1 & 0 & 0\\
2 & 1 & 0\\
6 & 1 & 1
\end{pmatrix}
\begin{pmatrix}
1 & -1 & 4\\
0 & 11 & 0\\
0 & 0 & -22
\end{pmatrix}"

Hence, the inverse of the matrix is "A = U^{-1} L^{-1}".

Since U and L are already in upper and lower triangular form, it is easy to find their inverses using Gauss-Jordan elimination.

Inverse of L:

"\left(\begin{array}{ccc|ccc} 
 1 & 0 & 0 & 1 & 0 & 0\\ 
 2 & 1 & 0 & 0 & 1 & 0\\
6 & 1 & 1 & 0 & 0 & 1
\end{array}\right) \sim (R_2 \rightarrow R_2- 2 R_1; R_3 \rightarrow R_3- 6 R_1) \sim"

"\left(\begin{array}{ccc|ccc} 
 1 & 0 & 0 & 1 & 0 & 0\\ 
 0 & 1 & 0 & -2 & 1 & 0\\
 0 & 1 & 1 & -6 & 0 & 1
\end{array}\right) \sim(R_3 \rightarrow R_3 - R_2) \sim
\left(\begin{array}{ccc|ccc} 
 1 & 0 & 0 & 1 & 0 & 0\\ 
 0 & 1 & 0 & -2 & 1 & 0\\
 0 & 0 & 1 & -4 & -1 & 1
\end{array}\right)"

Inverse of U:

"\left(\begin{array}{ccc|ccc} 
 1 & -1 & 4 & 1 & 0 & 0\\ 
 0 & 11 & 0 & 0 & 1 & 0\\
 0 & 1 & -22 & 0 & 0 & 1
\end{array}\right) \sim (R_2 \rightarrow \frac{R_2}{11}; R_3 \rightarrow \frac{R_3}{22}) \sim
\left(\begin{array}{ccc|ccc} 
 1 & -1 & 4 & 1 & 0 & 0\\ 
 0 & 1 & 0 & 0 & \frac{1}{11} & 0\\
 0 & 0 & 1 & 0 & 0 & -\frac{1}{22}
\end{array}\right) \sim (R_1 \rightarrow R_1 + R_2 - 4R_3) \sim
\left(\begin{array}{ccc|ccc} 
 1 & 0 & 0 & 1 & \frac{1}{11} & \frac{2}{11}\\ 
 0 & 1 & 0 & 0 & \frac{1}{11} & 0\\
 0 & 0 & 1 & 0 & 0 & -\frac{1}{22}
\end{array}\right)"

Hence, "A^{-1} = U^{-1}L^{-1} = 
\begin{pmatrix}
1 & \frac{1}{11} & \frac{2}{11}\\
0 & \frac{1}{11} & 0\\
0 & 0 & -\frac{1}{22}
\end{pmatrix}
\begin{pmatrix}
1 & 0 & 0\\
-2 & 1 & 0\\
-4 & -1 & 1
\end{pmatrix} = 
\begin{pmatrix}
\frac{1}{11} & -\frac{1}{11} & \frac{2}{11}\\
-\frac{2}{11} & \frac{1}{11} & 0 \\
\frac{2}{11} & \frac{1}{22} & -\frac{1}{22}
\end{pmatrix}".